Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Wu rolls two fair, six-sided dice. What is the probability

Expert replies
by swerve » Thu Mar 22, 2018 9:36 am
Wu rolls two fair, six-sided dice. What is the probability that Wu rolls at least one five but no sixes?

A. 5/36
B. 275/1296
C. 2/9
D. 1/4
E. 5/18

The OA is D.

Please, can any expert explain this PS question for me? I tried to solve it but I can't get the correct answer and I would like to know how to solve it in less than 2 minutes. I need your help. Thanks.
Join the discussion
Source: — Problem Solving |

Rolling at least one five and no sixes means:

one die with a 5 and the other with 1,2,3,4 OR

both dice with a 5

Because both of the above meet the criteria, you add their probabilities

Probability of a 5 on one die is 1/6. Probability of 1,2,3,4 on the other is 4/6 = 2/3. Multiplied together is 2/18.

Now because there are 2 ways for this to happen, the probability is 2x2/18 = 4/18 = 8/36

The probability of rolling two 5's is 1/6x1/6 = 1/36.

Adding 8/36 +1/36 = 9/36 = [spoiler]1/4, D[/spoiler]
Join the discussion

by [email protected] » Fri Mar 23, 2018 12:59 pm
Hi swerve,

We're told that Wu rolls two fair, six-sided dice. We're asked for the probability that Wu rolls AT LEAST one five but NO sixes. This question is all about 'probability math'; you might find it helpful to break the prompt into two smaller calculations:

Based on the types of outcomes that we are trying to achieve, there are two calculations to consider:

(1, 2, 3 or 4 on the 1st die)(5 on the 2nd die) = (4/6)(1/6) = 4/36
(5 on the 1st die)(1, 2, 3, 4 or 5 on the 2nd die) = (1/6)(5/6) = 5/36

Total Probability = 4/36 + 5/36 = 9/36 = 1/4

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Sun Jun 02, 2019 4:16 pm
swerve wrote:Wu rolls two fair, six-sided dice. What is the probability that Wu rolls at least one five but no sixes?

A. 5/36
B. 275/1296
C. 2/9
D. 1/4
E. 5/18
We can can break this up into scenarios:

Scenario 1:

P(5 on the first die and 1,2,3, or 4 on the second die)

1/6 x 4/6 = 4/36 = 1/9

Scenario 2:

P(1,2,3, or 4 on the first die and 5 on the second die)

4/6 x 1/6 = 4/36 = 1/9

Scenario 3:

P(5 on both dice)

1/6 x 1/6 = 1/36

Thus, the final probability is:

1/9 + 1/9 + 1/36 = 4/36 + 4/36 + 1/36 = 9/36 = 1/4.

Alternate Solution:

We see that of the 6 x 6 = 36 possible outcomes on the roll of two dice, only the outcomes of (5,1), (5,2), (5,3), (5,4); (1,5), (2,5), (3,5), (4,5) and (5,5) satisfy the required condition; a total of 9 outcomes. Thus, the probability that he rolls at least one 5 but no 6 is 9/36 = 1/4.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion